solve.
moolah has 24 coins that are all dimes and quarters.the value of the coins is $3.75. how many dimes and how many quarters does moolah have?
these are the equations i came up with
d=10 q=25 d+q=24 10d+25q=$3.75

Answer by mananth(16946)

d=10 q=25
d+q=24 .............................1
10d+25q=$3.75
3.75 = 375 cents
...
10d+25q=375.........................2
..
solve by addition
multiply (1) by -10
-10d-10q=-240
add it to (2)
25q-10q=375-240
15q=135
/15
q=135/15
q= 9
...
plug value of q in 1
d+q=24
d+9=24
d=15
..
CHECK
10*15+9*25
=150+225
=375.
m.ananth@hotmail.ca